A credit union offers 5.5% simple interest on a ceritificate of deposit (CD) and 3.5% simple interest on a savings account. If Mr. Sorkin invested $200 more in the CD than in the savings account, and the total interest after the first year was $245, how much was invested in each account?

I came up with 0.055(x + 200) + 0.035y=645 & (x + 200) + y = 1200, assuming that 1000 is needed to open a CD...

Do we presume that the intial invested in the CD is 1000 plus 200? Some people don't even know that most CD's is a minimum of a 1000 to open the account. I don't think so, let me know what you think when possible.

Thanks,
Alex Vargas

Answer by stanbon(75887)

credit union offers 5.5% simple interest on a ceritificate of deposit (CD) and 3.5% simple interest on a savings account. If Mr. Sorkin invested $200 more in the CD than in the savings account, and the total interest after the first year was $245, how much was invested in each account?
--------------------------------------------------
Equation:
interest + interest = 245 dollars
0.055(x+200) + 0.035x = 245
---
Multiply thru the equation by 1000 and solve for "x":
55x + 55*200 + 35x = 24500
90x = 13500
x = 150 (amt. invested in CD's)
---
x+200 = 350 (amt. invested in savings)
=======================================
Cheers,
Stan H.